A wire loop is in the x-y plane as shown. A curren
Answer #1

a) FL = Bo*I*L*sin(60) (towards -y axis)

FR = Bo*I*L*sin(60) (towards +y axis)

Ff = Bo*I*L*sin(90)

= Bo*I*L

Fbk = Bo*I*L*sin(90)

= Bo*I*L

b)


here FL and FR are same in magnitude but acts in the opposite direction.

Ff and Fbk are same in magnitude but acts in the opposite direction.


so, Net force = 0